# Coloring and density theorems for configurations of a given volume

### v2

@@ Abstract.
This is a treatise on finite point configurations spanning a fixed volume to be found in a single color-class of an arbitrary finite (measurable) coloring of the Euclidean space $\mathbb{R}^{n}$, or in a single large measurable subset $A\subseteq\mathbb{R}^{n}$. More specifically, we study vertex-sets of simplices, rectangular boxes, and parallelotopes, attempting to make progress on several open problems posed in the 1970s and the 1980s. As one of the highlights, we give the negative answer to a question of Erdős and Graham, by coloring the Euclidean plane $\mathbb{R}^{2}$ in $25$ colors without creating monochromatic rectangles of unit area. More generally, we construct a finite coloring of the Euclidean space $\mathbb{R}^{n}$ such that no color-class contains the $2^{m}$ vertices of any (possibly rotated) $m$-dimensional rectangular box of volume $1$. A positive result is still possible if rectangular boxes of merely sufficiently large volumes are sought in a single color-class of a finite measurable coloring of $\mathbb{R}^{n}$, and we establish it under an additional assumption $n\geqslant m+1$. Also, motivated by a question of Graham on reasonable bounds in his result on monochromatic axes-aligned right-angled $m$-dimensional simplices, we establish its measurable coloring and density variants with polylogarithmic bounds, again in dimensions $n\geqslant m+1$. Next, we generalize a result of Erdős and Mauldin, by constructing an infinite measure set $A\subseteq\mathbb{R}^{n}$ such that every $n$-parallelotope with vertices in $A$ has volume strictly smaller than $1$. Finally, some results complementing the literature on isometric embeddings of hypercube graphs and on the hyperbolic analogue of the Hadwiger–Nelson problem also follow as byproducts of our approaches.
@@ 1. Introduction
Euclidean Ramsey theory mainly investigates existence of a given point configuration in an arbitrary finite coloring of the Euclidean space $\mathbb{R}^{n}$ or its subsets. Its systematic study was initiated by Erdős, Graham, Montgomery, Rothschild, Spencer, and Straus [27, 28, 29] in the 1970s. On the other hand, geometric measure theory took a similar route in the 1980s and it started investigating point configurations that are found in arbitrary large measurable subsets of $\mathbb{R}^{n}$. Here we call a set $A\subseteq\mathbb{R}^{n}$ large if it has a positive density (as defined below) or, at least, a positive measure. Besides obvious connections to the Ramsey theory, the latter topic also shares the aesthetics with Szemerédi-type theorems from additive combinatorics. In fact, it can even be thought of as the arithmetic combinatorics in a geometric setting, typically formulated in the continuous parameter rather than a discrete one.
The present paper studies coloring and density theorems for rather flexible point configurations, from which we only require that they span a body of a given shape and a given volume. This particular subtopic has been initiated by Graham [34] and Erdős [25]. This treatise does not exhaust the subject, but it rather concentrates on the simplest configurations, such as simplices, rectangular boxes, and parallelotopes, as they exhibit different behaviors already. Our main goal is to show that many natural questions (or their variants) posed in the classical collections of open problems [26, 25] (also see [26, 12, 33]) are manageable by the techniques of harmonic analysis that have been developed in the meantime. However, quite a few basic questions are still left open.
Let us agree on the terminology. A finite coloring of $\mathbb{R}^{n}$ (or its subset $S$) is any partition of $\mathbb{R}^{n}$ (or $S$) into finitely many sets $\mathscr{C}_{1},\ldots,\mathscr{C}_{r}$, which are then called color-classes. A set $T$ is monochromatic if $T\subseteq\mathscr{C}_{j}$ for some index $1\leqslant j\leqslant r$. We say that the coloring is measurable if each of the color-classes $\mathscr{C}_{j}$ is a Lebesgue-measurable set. Similarly, we say that the coloring is Jordan-measurable if the boundary of $\mathscr{C}_{j}$ has Lebesgue measure 00 for every index $j$. Results that hold for an arbitrary finite (measurable) coloring of $\mathbb{R}^{n}$ (or an $S\subseteq\mathbb{R}^{n}$), perhaps with a prescribed number of colors, are sometimes referred to as coloring theorems.
Next, the Lebesgue measure of a measurable set $A\subseteq\mathbb{R}^{n}$ (i.e., the length, the area, or, generally, the volume) will simply be written as $|A|$. If $A$ is a measurable subset of the cube $[0,R]^{n}$ for some $R>0$, then we will refer to its density as the number $|A|/R^{n}$. On the other hand, for a general measurable $A\subseteq\mathbb{R}^{n}$ the relevant notion of largeness is the upper density, defined as
@@
for at least one index $1\leqslant j\leqslant r$. (An analogous claim holds for measurable colorings of the cube $[0,R]^{n}$.) Contrapositively, if we construct a measurable counterexample to a potential coloring result, then the corresponding density result is not possible either (cf. Theorems 3, 4, 6, and 8 below).
Admittedly, the measurability assumption is a seriously restricting one. There are many examples in the literature where the coloring result becomes significantly easier when only measurable colorings are considered; see Soifer’s book [50, Ch. 9]. For instance, the well-known Hadwiger–Nelson problem asks for a minimal number of colors needed to color $\mathbb{R}^{2}$ so that all pairs of points at distance $1$ apart receive different colors. It is easy to see that $7$ colors suffice. The exact minimal number has naturally been called the chromatic number of the plane. Back in 1981, Falconer [30] showed that the measurable chromatic number of $\mathbb{R}^{2}$ is at least $5$, but it took another $37$ years until de Grey [16] proved the same for the ordinary chromatic number. Also, from an old result by Furstenberg, Katznelson, and Weiss [32] it follows that a measurable set $A\subseteq\mathbb{R}^{2}$ satisfying $\bar{\textup{d}}_{2}(A)>0$ contains pairs of points at every sufficiently large distance from each other. This has been reproved independently by Falconer and Marstrand [31] and also by Bourgain [7]. As opposed to that, only recently Davies [14] solved a 1994 conjecture of Rosenfeld [50, §60.1] (popularized by Erdős), by showing that there exists a pair of points at an odd distance from each other for every (not necessarily measurable) finite coloring of $\mathbb{R}^{2}$. This was generalized to prime and polynomial distances by Davies, McCarty, and Pilipczuk [15]. A question by Kahle about distances of the form $n!$ [50, Problem 60.14] and the one by Soifer on distances of the form $2^{n}$ [50, Problem 60.16] remain open.
@@ 1.1. Triangles and simplices
Graham [34] answered positively a question of Gurevich, who asked whether, for positive integers $m\leqslant n$ and every finite coloring of $\mathbb{R}^{n}$, there exists an $m$-dimensional simplex of volume $1$ with all of its $m+1$ vertices colored the same [34, p. 89]. In fact, Graham could choose the simplices to be right-angled with axes-aligned perpendicular edges, i.e., the edges coming from a single vertex are parallel to the coordinate axes. Graham combined his observations with a general product theorem for colorings [34, Theorem 3], which he developed particularly for this purpose; also see the generalization that he proved jointly with Spencer [35]. His most general coloring theorem for simplices, namely [34, Theorem 4], reads as follows.
@@
By inspecting Graham’s proof (already for triangles) one notices that he first works on a huge part of the integer lattice $\mathbb{Z}^{n}$ [34, Theorem 1], the size of which is an iterated van der Waerden number in the number of colors used. Then he scales the lattice to obtain the result in $\mathbb{R}^{n}$, but also wonders [34, p. 97] if a proof using only a reasonable number of integer points (or, strictly speaking, only reasonably large integer-point triangles) is possible. A natural continuous-parameter version of Graham’s question is the one that considers only colorings of a sufficiently large cube $[0,R]^{n}\subseteq\mathbb{R}^{n}$.
@@
Let $m,n$ be fixed positive integers such that $n\geqslant m$. Is there a reasonable lower bound (i.e., not of the Ackermann type) on the number $R=R(r)\in(0,\infty)$ such that, for every coloring of the $n$-cube $[0,R]^{n}$ in $r$ colors, there exists a right-angled $m$-dimensional simplex of unit volume with monochromatic vertices?
@@
    If $R\in(0,\infty)$ and if the $n$-cube $[0,R]^{n}$ is measurably colored in $r$ colors, then
@@
In both parts of the theorem, the $m$-simplex can be chosen such that $m-1$ of its edges from the right-angled vertex are parallel to the coordinate vectors $\mathbbm{e}_{1},\ldots,\mathbbm{e}_{m-1}$, while the remaining edge is parallel to the linear span of $\mathbbm{e}_{m},\ldots,\mathbbm{e}_{n}$.
@@
Graham suggested (without sharing many details) the possibility of a density variant of his discrete result [34, Theorem 1] in Concluding Remarks [34, p. 96], but he mentioned general triangles only, i.e., not axes-aligned or even right-angled ones, which makes the problem easier. Indeed, the corresponding analogue in the plane $\mathbb{R}^{2}$ is indeed quite easy: it is an old observation of Erdős, first published as an exercise in the popular magazine Matematikai Lapok; see the comments in [24, p. 122].
@@
The proof is a simple consequence of the one-dimensional Steinhaus theorem [53] and it was elaborated on in [12, p. 182]. The corresponding problem for right-angled triangles Erdős mentioned as being open [25, p. 324].
Problem 2 (Erdős [25]).
Let $A\subseteq\mathbb{R}^{2}$ be a measurable set with infinite area. Must $A$ contain the vertices of a right triangle of area $1$?
In the last problem we cannot also ask from the triangle to be axes-aligned. Indeed, the example of a rotated thin infinite strip/tube shows that, if $A\subseteq\mathbb{R}^{n}$ merely has infinite volume, then the perpendicular sides of right-angled unit triangles/simplices with vertices in $A$ can avoid any prescribed tuple of orthogonal directions. The same example also shows that the simplices with vertices in a set $A$ of merely positive area can all have disproportional ratios of their edge-lengths. For this reason we find non-trivial the following consequence of Theorem 1(a), which deals with sets of positive upper Banach density.
@@
Even a positive solution to Problem 2 and its higher-dimensional generalizations still would not cover the quantitative aspect of Corollary 2.
An interesting related problem is an old question of Erdős on triangles in planar sets $A$ of merely sufficiently large finite area [24, p. 122–123].
@@
Problem 3 has since been featured in numerous collections of open problems, such as the book Unsolved problems in geometry by Croft, Falconer, and Guy [12, Problem G13], Bloom’s website Erdős problems [6, #352], or the expository paper by Mauldin [47]. The techniques in this paper are insufficient to detect configurations in sets that are only known to have large measure.
@@ 1.2. Rectangles and rectangular boxes
(…) can be generalized to configurations which are not simplexes. [34, p. 96].
@@
The question is: Is this also true for rectangles? Or perhaps parallelograms? [26, p. 15].
To the best of author’s knowledge, these problems, either for rectangles or for parallelograms, have not been addressed in the literature so far. They were explicitly mentioned as being open in the papers by Adhikari and Rath [2, 3] and in Mathematical Reviews MR0558877 by K. Steffens and MR2106573 by S. Oates-Williams. They were also recently stated among the open questions in the 2015 edition of the book Rudiments of Ramsey theory by Graham and Butler [33, p. 56].
@@
Note that the rectangles can be arbitrary rotated. We are able to show that Problem 4 has the negative answer.
@@
In both parts of the theorem, rectangular boxes can be chosen to have $m-1$ edges parallel to the first $m-1$ coordinate vectors and the last edge parallel to the linear span of the remaining $n-m+1$ coordinate vectors.
Theorem 5 is only new when the ambient dimension $n$ satisfies $m+1\leqslant n\leqslant 2m-1$. It follows from the existing literature in the cases $n\geqslant 2m$, as Lyall and Magyar [46, Theorem 1.1(i)] showed that a set $A\subseteq\mathbb{R}^{2m}$ of positive upper Banach density contains the vertices of an $m$-cube with every sufficiently large edge-length. The critical case $n=m\geqslant 2$ is not covered by our theorem and it remains open. The analogous result for $n=m=1$ clearly fails, as is seen by taking $A=(-1/5,1/5)+\mathbb{Z}$. Thus, the following problem remains.
@@
It is also unclear if the measurability condition can be removed from the statement of Theorem 5(b).

### v3

@@ Abstract.
This is a treatise on finite point configurations spanning a fixed volume to be found in a single color-class of an arbitrary finite (measurable) coloring of the Euclidean space $\mathbb{R}^{n}$, or in a single large measurable subset $A\subseteq\mathbb{R}^{n}$. More specifically, we study vertex-sets of simplices, rectangular boxes, and parallelotopes, attempting to make progress on several open problems posed in the 1970s and the 1980s. As one of the highlights, we give a negative answer to a question of Erdős and Graham, by coloring the Euclidean plane $\mathbb{R}^{2}$ in $25$ colors without creating monochromatic rectangles of unit area. More generally, we construct a finite coloring of the Euclidean space $\mathbb{R}^{n}$ such that no color-class contains the $2^{m}$ vertices of any (possibly rotated) $m$-dimensional rectangular box of volume $1$. A positive result is still possible if rectangular boxes of merely sufficiently large volumes are sought in a single color-class of a finite measurable coloring of $\mathbb{R}^{n}$, and we establish it under an additional assumption $n\geqslant m+1$. Also, motivated by a question of Graham on reasonable bounds in his result on monochromatic axes-aligned right-angled $m$-dimensional simplices, we establish its measurable coloring and density variants with polylogarithmic bounds, again in dimensions $n\geqslant m+1$. Next, we generalize a result of Erdős and Mauldin, by constructing an infinite measure set $A\subseteq\mathbb{R}^{n}$ such that every $n$-parallelotope with vertices in $A$ has volume strictly smaller than $1$. Finally, some results complementing the literature on isometric embeddings of hypercube graphs and on the hyperbolic analogue of the Hadwiger–Nelson problem also follow as byproducts of our approaches.
@@ 1. Introduction
Euclidean Ramsey theory mainly investigates the existence of a given point configuration in an arbitrary finite coloring of the Euclidean space $\mathbb{R}^{n}$ or its subsets. Its systematic study was initiated by Erdős, Graham, Montgomery, Rothschild, Spencer, and Straus [28, 29, 30] in the 1970s. On the other hand, geometric measure theory took a similar route in the 1980s and it started investigating point configurations that are found in arbitrary large measurable subsets of $\mathbb{R}^{n}$. Here we call a set $A\subseteq\mathbb{R}^{n}$ large if it has a positive density (as defined below) or, at least, a positive measure. Besides obvious connections to the Ramsey theory, the latter topic also shares the aesthetics with Szemerédi-type theorems from additive combinatorics. In fact, it can even be thought of as arithmetic combinatorics in a geometric setting, typically formulated in a continuous parameter rather than a discrete one.
The present paper studies coloring and density theorems for rather flexible point configurations, from which we only require that they span a body of a given shape and a given volume. This particular subtopic has been initiated by Graham [36] and Erdős [26]. This treatise does not exhaust the subject, but rather concentrates on the simplest configurations, such as simplices, rectangular boxes, and parallelotopes, as they exhibit different behaviors already. Our main goal is to show that many natural questions (or their variants) posed in the classical collections of open problems [27, 26] (also see [12, 35]) are manageable by the techniques of harmonic analysis that have been developed in the meantime. However, quite a few basic questions are still left open.
Let us agree on the terminology. A finite coloring of $\mathbb{R}^{n}$ (or its subset $S$) is any partition of $\mathbb{R}^{n}$ (or $S$) into finitely many sets $\mathscr{C}_{1},\ldots,\mathscr{C}_{r}$, which are then called color-classes. A set $T$ is monochromatic if $T\subseteq\mathscr{C}_{j}$ for some index $1\leqslant j\leqslant r$. We say that the coloring is measurable if each of the color-classes $\mathscr{C}_{j}$ is a Lebesgue-measurable set. Similarly, we say that the coloring is Jordan-measurable if the boundary of $\mathscr{C}_{j}$ has Lebesgue measure 0 for every index $j$. Results that hold for an arbitrary finite (measurable) coloring of $\mathbb{R}^{n}$ (or an $S\subseteq\mathbb{R}^{n}$), perhaps with a prescribed number of colors, are sometimes referred to as coloring theorems.
Next, the Lebesgue measure of a measurable set $A\subseteq\mathbb{R}^{n}$ (i.e., the length, the area, or, generally, the volume) will simply be written as $|A|$. If $A$ is a measurable subset of the cube $0,R]^{n}$ for some $R>0$, then we will refer to its density as the number $|A|/R^{n}$. On the other hand, for a general measurable $A\subseteq\mathbb{R}^{n}$ the relevant notion of largeness is the upper density, defined as
@@
for at least one index $1\leqslant j\leqslant r$. (An analogous claim holds for measurable colorings of the cube $0,R]^{n}$.) Contrapositively, if we construct a measurable counterexample to a potential coloring result, then the corresponding density result is not possible either (cf. Theorems 3, 4, 6, and 8 below).
Admittedly, the measurability assumption is a seriously restricting one. There are many examples in the literature where the coloring result becomes significantly easier when only measurable colorings are considered; see Soifer’s book [55, Ch. 9]. For instance, the well-known Hadwiger–Nelson problem asks for a minimal number of colors needed to color $\mathbb{R}^{2}$ so that all pairs of points at distance $1$ apart receive different colors. It is easy to see that $7$ colors suffice. The exact minimal number has naturally been called the chromatic number of the plane. Back in 1981, Falconer [31] showed that the measurable chromatic number of $\mathbb{R}^{2}$ is at least $5$, but it took another $37$ years until de Grey [16] proved the same for the ordinary chromatic number. Also, from an old result by Furstenberg, Katznelson, and Weiss [34] it follows that a measurable set $A\subseteq\mathbb{R}^{2}$ satisfying $\bar{\textup{d}}_{2}(A)>0$ contains pairs of points at every sufficiently large distance from each other. This has been reproved independently by Falconer and Marstrand [32] and also by Bourgain [7]. As opposed to that, only recently Davies [13] solved a 1994 conjecture of Rosenfeld [55, §60.1] (popularized by Erdős), by showing that there exists a pair of points at an odd distance from each other for every (not necessarily measurable) finite coloring of $\mathbb{R}^{2}$. This was generalized to prime and polynomial distances by Davies, McCarty, and Pilipczuk [15]. A question by Kahle about distances of the form $n!$ [55, Problem 60.14] and the one by Soifer on distances of the form $2^{n}$ [55, Problem 60.16] remain open.
@@ 1.1. Triangles and simplices
Graham [36] answered positively a question of Gurevich, who asked whether, for positive integers $2\leqslant m\leqslant n$ and every finite coloring of $\mathbb{R}^{n}$, there exists an $m$-dimensional simplex of volume 11 with all of its $m+1$ vertices colored the same [36, p. 89]. In fact, Graham could choose the simplices to be right-angled with axes-aligned perpendicular edges, i.e., the edges coming from a single vertex are parallel to the coordinate axes. Graham combined his observations with a general product theorem for colorings [36, Theorem 3], which he developed particularly for this purpose; also see the generalization that he proved jointly with Spencer [37]. His most general coloring theorem for simplices, namely [36, Theorem 4], reads as follows.
@@
By inspecting Graham’s proof (already for triangles) one notices that he first works on a huge part of the integer lattice $\mathbb{Z}^{n}$ [36, Theorem 1], the size of which is an iterated van der Waerden number in the number of colors used. Then he scales the lattice to obtain the result in $\mathbb{R}^{n}$, but also wonders [36, p. 97] if a proof using only a reasonable number of integer points (or, strictly speaking, only reasonably large integer-point triangles) is possible. A natural continuous-parameter version of Graham’s question is the one that considers only colorings of a sufficiently large cube $0,R]^{n}\subseteq\mathbb{R}^{n}$.
@@
Let $m,n$ be fixed positive integers such that $n\geqslant m\geqslant 2$. Is there a reasonable lower bound (i.e., not of the Ackermann type) on the number $R=R(r)\in(0,\infty)$ such that, for every coloring of the $n$-cube $0,R]^{n}$ in $r$ colors, there exists a right-angled $m$-dimensional simplex of unit volume with monochromatic vertices?
@@
    If $R\in(0,\infty)$ and if the $n$-cube $0,R]^{n}$ is measurably colored in $r$ colors, then
@@
In both parts of the theorem, the $m$-simplex can be chosen such that m1m-1 of its edges from the right-angled vertex are parallel to the coordinate vectors $\mathbbm{e}_{1},\ldots,\mathbbm{e}_{m-1}$, while the remaining edge is parallel to the linear span of $\mathbbm{e}_{m},\ldots,\mathbbm{e}_{n}$.
@@
Graham suggested (without sharing many details) the possibility of a density variant of his discrete result [36, Theorem 1] in Concluding Remarks [36, p. 96], but he mentioned general triangles only, i.e., not axes-aligned or even right-angled ones, which makes the problem easier. Indeed, the corresponding analogue in the plane $\mathbb{R}^{2}$ is, however, quite easy: it is an old observation of Erdős, first published as an exercise in the popular magazine Matematikai Lapok; see the comments in [24, p. 122].
@@
The proof is a simple consequence of the one-dimensional Steinhaus theorem [58] and it was elaborated on in [12, p. 182]. The corresponding problem for right-angled triangles was mentioned by Erdős [26, p. 324] and attributed to him and Mauldin.
Problem 2 (Erdős and Mauldin [26]).
Let $A\subseteq\mathbb{R}^{2}$ be a measurable set with infinite area. Must $A$ contain the vertices of a right triangle of area $1$? ✓
Very recently Koizumi [40] answered this question positively, claiming to have been motivated by the ideas from a recent paper by Predojević and the present author [46]. The techniques from [46, 40] rely on the Lebesgue density theorem and thus cannot recover any of the inherently quantitative results of the present paper. Another four questions by Erdős and Mauldin (on other shapes) similar to Problem 2 can be found in [26, p. 323–324] or gathered under a single problem on Bloom’s website Erdős problems [6, #353]. They were all answered either in [46] or in [40].
In the last problem we cannot also ask from the triangle to be axes-aligned. Indeed, the example of a rotated thin infinite strip/tube shows that, if $A\subseteq\mathbb{R}^{n}$ merely has infinite volume, then the perpendicular sides of right-angled unit triangles/simplices with vertices in $A$ can avoid any prescribed tuple of orthogonal directions. The same example also shows that the simplices with vertices in a set $A$ of merely positive area can all have disproportional ratios of their edge-lengths. For this reason we find non-trivial the following consequence of Theorem 1(a), which is quantitative and it deals with sets of positive upper Banach density.
@@
An interesting related problem is an old question of Erdős on triangles in planar sets $A$ of merely sufficiently large finite area.
@@
It seems that Erdős asked this question first in [24, p. 122–123], and then repeated it in [25, p. 30] and [26, p. 323]. Problem 3 has since been featured in numerous collections of open problems, such as the book Unsolved problems in geometry by Croft, Falconer, and Guy [12, Problem G13], the website Erdős problems [6, #352], or the expository paper by Mauldin [51]. The techniques in this paper are insufficient to detect configurations in “scattered” sets that are only known to have large measure.
@@ 1.2. Rectangles and rectangular boxes
(…) can be generalized to configurations which are not simplexes. [36, p. 96].
@@
The question is: Is this also true for rectangles? Or perhaps parallelograms? [27, p. 15].
To the best of the author’s knowledge, these problems, either for rectangles or for parallelograms, have not been addressed in the literature so far. They were explicitly mentioned as being open in the papers by Adhikari and Rath [2, 3] and in Mathematical Reviews MR0558877 by K. Steffens and MR2106573 by S. Oates-Williams. They were also recently stated among the open questions in the 2015 edition of the book Rudiments of Ramsey theory by Graham and Butler [35, p. 56].
@@
Note that the rectangles can be arbitrarily rotated. We are able to show that Problem 4 has a negative answer.
@@
The minimal number of colors enabling a coloring of $\mathbb{R}^{n}$ with the property from Theorem 4 grows at least exponentially in $n$. This is seen by taking $m=1$ and consulting the literature on the higher-dimensional Hadwiger–Nelson problem, a.k.a. the chromatic number of $\mathbb{R}^{n}$; see [33]. On the other hand, we will see that the number of colors needed in the proof presented in Section 3 grows superexponentially in $n$. It would be interesting to obtain reasonably sharp bounds on the cardinality of the minimal coloring.
@@
In both parts of the theorem, rectangular boxes can be chosen to have m1m-1 edges parallel to the first m1m-1 coordinate vectors and the last edge parallel to the linear span of the remaining nm+1n-m+1 coordinate vectors.
Theorem 5 is only new when the ambient dimension $n$ satisfies $m+1\leqslant n\leqslant 2m-1$. It follows from the existing literature in the cases $n\geqslant 2m$, as Lyall and Magyar [50, Theorem 1.1(i)] showed that a set $A\subseteq\mathbb{R}^{2m}$ of positive upper Banach density contains the vertices of an $m$-cube with every sufficiently large edge-length. The critical case $n=m\geqslant 2$ is not covered by our theorem and it remains open. The analogous result for $n=m=1$ clearly fails, as is seen by taking $A=(-1/5,1/5)+\mathbb{Z}$. Thus, the following problem remains.
@@
It is hard to say if Problem 5 is within the reach of harmonic analysis techniques, but it is expected to be difficult. Namely, when one sets up the harmonic analysis “machinery” it seems that the relevant analytical objects are just slightly “too singular” to be handled using the currently known ideas. Also, its case $n=2$ is formally similar to Problem 7 below, for triangles in the plane, which is open as well. It is also unclear if the measurability condition can be removed from the statement of Theorem 5(b).